brooklynjojo6:

can someone help me bro my teacher gave me this question for a 100 on 5 assignments x² - 5x + 6 = 0.

umm:

What do you need help with? Factoring, quadratic formula, completing the square, or finding the discriminant?

somnium:

if youre looking for the root x values, theyre x=2 and x=3. really the easiest way to solve it is to throw it onto a desmos graph, but you could use quadratic equation to solve it if desired

Aratox:

To solve the quadratic equation \(x^2 - 5x + 6 = 0\), we can factor it or use the quadratic formula. Factoring involves finding two numbers that multiply to \(6\) and add to \(-5\), which are \(2\) and \(3\). So, we rewrite the equation as \((x - 2)(x - 3) = 0\). Then, by the zero product property, we set each factor equal to zero to find the solutions: \(x - 2 = 0\) gives \(x = 2\) and \(x - 3 = 0\) gives \(x = 3\).